### v4

# Microwave Signatures of Topological Superconductivity in Planar Josephson Junctions

Abstract
Planar Josephson junctions provide a platform to host topological superconductivity which, through manipulating Majorana bound states (MBS), could enable fault-tolerant quantum computing. However, what constitutes experimental signatures of topological superconductivity and how MBS can be detected remains strongly debated. In addition to spurious effects that mimic MBS, there is a challenge to discern the inherent topological signals in realistic systems with many topologically-trivial Andreev bound states, determining the transport properties of Josephson junctions. Guided by the advances in microwave spectroscopy, we theoretically study Al/InAs-based planar Josephson junction embedded into a radio-frequency superconducting quantum interference device to identify microwave signatures of topological superconductivity. Remarkably, by exploring the closing and reopening of a topological gap, we show that even in a wide planar Josephson junction with many Andreev bound states, such a topological signature is distinguishable in the resonance frequency shift of a microwave drive and the “half-slope” feature of the microwave absorption spectrum. Our findings provide an important step towards experimental detection of non-Abelian statistics and implementing scalable topological quantum computing.
The interest in topological superconductors reflects their fascinating fundamental properties and potential applications using manipulation of Majorana bound states (MBS) Kitaev (2003); Nayak et al. (2008); Alicea et al. (2011); Sarma et al. (2015); Matos-Abiague et al. (2017); Güngördü and Kovalev (2022); Laubscher and Klinovaja (2021). However, detecting such topological superconductivity and MBS remains a long-standing challenge Amundsen et al. (2024); Yu et al. (2021); Chen et al. (2019); Lee et al. (2012). Revisiting Sr2RuO4, expected for decades to be $p$-wave spin-triplet which could support MBS Mackenzie and Maeno (2003); Rice (2004); Nelson et al. (2004); Das Sarma et al. (2006), or quantized zero-bias conductance peak (ZBCP) Sengupta et al. (2001) as an experimental MBS signature Mourik et al. (2012) both provide cautionary guidance for various difficulties in detection of superconducting properties Žutić and Mazin (2005); Chen et al. (2019); Das Sarma and Pan (2021).
We focus on planar Josephson junctions (JJs) Shabani et al. (2016); Hell et al. (2017); Pientka et al. (2017); Fornieri et al. (2019); Ren et al. (2019); Setiawan et al. (2019); Dartiailh et al. (2021a); Banerjee et al. (2023), where topological superconductivity is realized through proximity effects in the normal region (N) of two-dimensional electron gas (2DEG), which separates superconducting (S) regions, shown in Fig. 1(a). Such JJs, with a phase or gate control and multi-terminal geometries suitable for probing non-Abelian statistics Fu and Kane (2008); Zhou et al. (2020, 2022), could overcome limitations for topological superconductivity in 1D nanowires requiring fine-tuned parameters Laubscher and Klinovaja (2021); Lutchyn et al. (2010); Oreg et al. (2010) and MBS detection through ZBCP.
Our approach builds on advances in circuit quantum electrodynamics and highly-sensitive microwave spectroscopy of Andreev bound states (ABS) Chtchelkatchev and Nazarov (2003); Despósito and Levy Yeyati (2001); Bretheau et al. (2013a); Reynoso et al. (2012); Tosi et al. (2019). Due to electron-hole symmetry, ABS come in pairs and with energies ±$\pm E_{\mathrm{ABS}}(\phi)$, that depend on the relative superconducting phase, $\phi$, between the two S regions. In contrast, MBS are topologically-protected ABS with $E\approx 0$ localized to the ends of the junction that cannot be doubly occupied. To investigate these states, we theoretically consider an experimentally accurate model system where the planar JJ is embedded into a radio frequency superconducting quantum interference device (rf-SQUID), driven at a microwave frequency, $f_{d}$. The absorbed photons with $E=hf_{d}$ cause transitions between the ABS and change the JJ’s critical current, $I_{c}$. By coupling the rf-SQUID to a microwave resonator, one can calculate this effect from the resulting change in the resonant frequency, $f_{r}$, of the resonator via a capacitively coupled transmission line Park et al. (2020); Pozar (2012); Elfeky et al. (2023a), as shown in Fig. 1. This approach, using numerical discretization, allows us to calculate the ABS spectrum and topological phase diagram of the planar JJ, as well as the microwave spectra and the shift in the critical current.
Figure 1: (a) A planar JJ formed by two Al superconducting (S) regions covering InAs-based 2DEG normal region (N). Coordinate axes and JJ dimensions are indicated. Alternating current, $I$, flows along ±$\pm\hat{x}$; the applied magnetic field is ${\bm{B}}$. (b)-(d) Optical images of an example fabricated SQUID resonator, which we consider in our theoretical results. Gold: Al, dark brown: the etched mesa, and light brown: gate and a flux line. (c) Device diagram with resonator capacitively coupled to a transmission line and inductively to a superconducting loop with a flux $\Phi$ and a single JJ. (d) Schematic description from (c) of the JJ, SQUID, microwave drive line inducing fluctuations in $\Phi\rightarrow\Phi+\delta\Phi(t)$, resonator, and transmission line with input (output) power $P_{\text{in}}$ ($P_{\text{out}}$).
Realistic planar JJs, in contrast with many microwave studies focusing on a few ABS which is a characteristic for nanowires, atomic contacts, or constrictions, can have hundreds of topologically-trivial ABS, which complicate detecting inherent properties of topological superconductivity and MBS Deacon et al. (2010); Zhang et al. (2022); Janvier et al. (2015); Wallraff et al. (2004); Tosi et al. (2019); van Woerkom et al. (2017); Hays et al. (2021); Pekker et al. (2013); Matute-Cañadas et al. (2022). Based on the experimental parameters of our fabricated JJs, we theoretically identify microwave signatures for closing and reopening of a topological gap. Our calculations show that such signatures are distinguishable in the transitions between the ABS due to a microwave drive of frequency, $f_{d}$, which change the current-phase relation (CPR), $I(\phi)=\Sigma_{n}I_{n}^{\mathrm{ABS}}$ that depends on the occupation of the ABS levels. This change in the CPR results in a shift in $f_{r}$, measured in the microwave circuit Park et al. (2020); Pozar (2012); Elfeky et al. (2023a).
In microwave detection, a quality factor $\gtrsim 10,000$ and a linewidth of $0.1-1\;$MHz yield fast ($\sim\mu$s) and neV energy resolution of ABS Tosi et al. (2019); Hays et al. (2018). In contrast to ZBCP measurements, this high sensitivity and fermion-parity conservation, which enables the probing of quasiparticle poisoning and dynamical properties of superconducting devices Kurilovich et al. (2021); Kos et al. (2013); Hays et al. (2018); Elfeky et al. (2023a), make our findings important beyond signatures of topological superconductivity and MBS. These findings could also advance the microwave studies in many other systems with a large number of ABS, for example, in superconducting spintronics Linder and Robinson (2015); Eschrig (2015); Cai et al. (2023); Amundsen et al. (2024), in nonreciprocal phenomena such as diode effects Amundsen et al. (2024); Baumgartner et al. (2022); Dartiailh et al. (2021a); Pekerten et al. (2022); Lotfizadeh et al. (2024), or in emerging superconducting qubits Krantz et al. (2019); Tafuri (2019) (ed.); Pita-Vidal et al. (2020).
We consider a model Al/InAs-based JJ with $L=4\;\mu$m and $W_{N}=80\;$nm [see Fig. 1], embedded in an rf-SQUID, that is inductively coupled to a hanger $\lambda/4$ coplanar waveguide resonator for our calculations. The model resonator is capacitively coupled to a transmission feed line, allowing for the complex transmission and the resonant response to be measured. We use $f_{r}\sim 7\;$GHz, estimated from experimentally relevant circuit parameters Park et al. (2020); Pozar (2012). We therefore present our numerical results with $f_{d}$ up to a few $f_{r}$ ($\sim 40$ GHz), unless showing the full range of $h\,f_{d}\lesssim 2\Delta_{0}\sim 120$ GHz is instructive.
Our model for the microwave-driven JJ is described by the Bogoliubov-de Gennes (BdG) Hamiltonian Pientka et al. (2017); Zhou et al. (2022)
$H_{0}$ $=\left[\frac{\mathbf{p}^{2}}{2m^{\ast}}-\mu\left(x,y\right)+\frac{\alpha}{\hbar}\left(p_{x}\sigma_{y}-p_{y}\sigma_{x}\right)\right]\tau_{z}$
$\quad-\frac{g^{\ast}\mu_{B}}{2}\mathbf{B}\cdot\bm{\sigma}+\Delta\left(x,y\right)\tau_{+}+\Delta^{\ast}\left(x,y\right)\tau_{-}\;,$ (1)
for which numerically solve the corresponding discretized eigenvalue problem SM_ to obtain ABS. Here, ${\bf p}$ is the momentum, $\mu\left(x,y\right)$ is the chemical potential, $\alpha$ is the Rashba spin-orbit coupling (SOC) strength, $\mathbf{B}=B\,\hat{y}$ is the in-plane magnetic field and $\mu_{B}$ is the Bohr magneton, while $m^{\ast}$ and $g^{\ast}$ are the effective mass and $g$-factor, respectively. The proximity-induced superconductivity in the InAs 2DEG is described by the pair potential $\Delta(x,y)=\Delta_{0}\,\Theta(|x-W_{N}/2|)\,\exp{(i\,\phi(x,y)/2})$, where $\Theta$ is the step function and in $\phi(x,y)=\phi_{0}\,\mathrm{sgn}{(x)}$, $\phi_{0}$ is uniform in each S region. Similarly, we take $\Delta_{0}\rightarrow\Delta_{0}(B,T)$ to be real and uniform in the S regions, having the usual BCS dependence on $B$ and temperature $T$ Tinkham (1996); Zhou et al. (2022). $\sigma_{i}$ ($\tau_{i}$) are Pauli (Nambu) matrices in the spin (particle-hole) space, and $\tau_{\pm}=(\tau_{x}\pm i\tau_{y})/2.$
Figure 2: (a), (c) ABS spectra as a function of $\phi$ for JJ from Fig. 1(a). $W_{S}=600\;$nm, $W_{N}=80\;$nm, $L=4\;\mu$m, and $B=0.2\;$T. (b), (d) The corresponding topological charge, $Q_{D}$, as a function of $B$ and $\phi$. (a), (b) For $\mu=0.1\;$meV, the gapped ABS are trivial for all $\phi$. (c) For $\mu=1.0\;$meV, there is a gap closing and reopening. The inset: MBS at $E\approx 0$ (red solid line). (d) A topological transition as $\phi$ is varied at $B=0.2\;$T, corresponding to the gap closing points in (c).
Figure 3: (a), (b) Transition lines $\widetilde{\mathcal{T}}(f,\phi)$, calculated using Eq. (4), corresponding to $hf_{d}=\Delta E=|E_{n}-E_{m}|$ between ABS from Figs. 2(a) and (c), with a Lorentzian broadening parameter $\epsilon=30\;$mK. In the schematic of pair (c) and single-particle (d) transitions where the ground state is denoted by a blue line, the sum of the arrow lengths in (c) gives $\Delta E$. Resolving (e) pair and (f) single-particle transitions from (a) up to 404040\;40GHz shows no gap closing and opening. In contrast a similarly resolved (g) pair and (h) single-particle transitions from (b) reveal signatures of topological superconductivity. The inset: an enlarged part of the Fig. 2(c) spectra. The transitions in the circles “A” and “B” of the inset lead to the V-shaped features “A” and “B” in (g), signaling a gap closing and reopening. Transitions in the circle “C” lead to the V-shaped feature “C” in (h).
We choose material parameters consistent with our fabricated epitaxial Al/InAs JJs which support robust proximity-induced superconductivity and topological transitions Dartiailh et al. (2021a): $\Delta_{0}=0.23\;$meV, $m^{\ast}=0.027m_{0}$ where $m_{0}$ is the electron mass, $g^{\ast}=10$ for InAs, $\alpha=10$ meVnm, critical magnetic field $B_{c}(T=0)=1.6\;$T and temperature $T=30\;$mK. The corresponding ABS spectra are shown in Fig. 2. In the top row, for $\mu=0.1\;$meV and $B=0.2\;$T, the JJ is in the topologically trivial phase for all $\phi$. In Fig. 2(a) there is no gap closing at $B=0.2\;$T, while in Fig. 2(b) the calculated topological charge Nayak et al. (2008), $Q_{D}=1$, confirms this trivial phase.The mirror symmetry in the system with respect to the middle of the junction along the $x$-direction allows for an effective time reversal symmetry, leading to a system that can support MBS in a wider range of $\phi$ values around $\phi=\pi$, as seen in Figure 2(b) and (d) Pientka et al. (2017). For a finite $W_{S}$, there are normal reflections and an imperfect transparency at the N/S interface, causing the topological phase diagram to deviate from the typical “diamond shape” Pientka et al. (2017); Pekerten et al. (2022, 2024) with a trivial phase even at $\phi=\pi$ and small $B$. In contrast, for a larger $\mu$, the JJ in our model undergoes topological transitions even for $B\sim 0$ at $\phi=\pi$, leading to a wider range of $\phi$ that can support MBS for higher $B$ [Fig. 2(d)]. At $B=0.2\;$T, MBS near $E\sim 0$ is seen in the spectrum in Fig. 2(c) and the inset for $0.8\pi<\phi<1.2\pi$2 2 In Supplemental Material (SM) SM_ , we complement these results with CPR, topological gap, the MBS localization, as well as topological phase transitions for larger $\mu$ as $B$ is varied.
We next discuss the signatures of the topological gap closing and reopening in Fig. 2(c) using the microwave spectroscopy with the setup from Figs. 1(b)-(d). The microwave drive induces a flux fluctuation, $\Phi(t)=\Phi(0)+\delta\Phi(t)$, in the rf-SQUID around its unperturbed value $\Phi(0)$, which leads to the JJ phase fluctuation $\delta\phi(t)$, described by the perturbed Hamiltonian Despósito and Levy Yeyati (2001); Zazunov et al. (2003); Janvier et al. (2015); Hays et al. (2021); Olivares et al. (2014); Väyrynen et al. (2015)
$H(\phi+\delta\phi(t))$ $=H_{0}(\phi)+\delta\phi(t)\,H_{\mathrm{MW}},$ (2)
where the perturbation $H_{\mathrm{MW}}=\partial H_{0}(\phi)\,/\,\partial\phi$ is given by
$H_{\mathrm{MW}}=-i\,(\Delta_{0}/2)\,\left(e^{-i\phi(\mathbf{r})/2}\,\tau_{+}-e^{i\phi(\mathbf{r})/2}\,\tau_{-}\right).$ (3)
For our model experimental setup based on fabricated JJs embedded in microwave resonator circuits, we can examine the validity of this approach by a simple estimate, $\delta\phi/2\pi\sim M\,I_{\mathrm{MW}}/\Phi_{0}<M\,I_{c}/\Phi_{0}\approx 0.05$, where we use an experimentaly relevant value of $M\sim 120\;$pH for the mutual inductance and a critical current of $I_{c}\lesssim 1\;\mu$A as an upper bound for the microwave driving current $I_{\mathrm{MW}}$ Dartiailh et al. (2021b); Phan et al. (2023), thus justifying the treatment of the effects of the microwave drive as a perturbation Väyrynen et al. (2015); Wesdorp et al. (2024); Chidambaram et al. (2022). The transitions caused by this perturbation conserve the fermion parity Bretheau et al. (2017); Janvier et al. (2015); Kos et al. (2013); Väyrynen et al. (2015); SM_ , which reduces the number of the allowed transition lines shown in Figs. 3(a) and (b). We can then distinguish the pair transitions, between a Cooper pair and two ABS, or the single-particle transitions between two ABS, as denoted in Figs. 3(c) and (d). In the BdG Hamiltonian, pair (single-particle) transitions correspond to transitions between the states with opposite (same) energy sign.
With high energy- and time-resolution, microwave spectroscopy provides information about the allowed transitions lines, $h\,f_{d}^{nm}=|E_{m}-E_{n}|$, with the transition probability $|\mathcal{M}_{nm}|^{2}$ governed by the matrix element $\mathcal{M}_{nm}=\braket{\psi_{m}}{H_{\mathrm{MW}}}{\psi_{n}}$, where $\psi_{n}$ is the eigenstate with eigenenergy $E_{n}$ of $H_{0}$ [Eq. (1)]. Our calculated results for planar JJs in Fig. 3 show the allowed transition lines
$\mathcal{T}(f_{d},\phi)=\sum_{mn}|\mathcal{M}_{nm}|^{2}\mathcal{F}(E_{m},E_{n})\,\mathcal{L}(E_{mn}-hf_{d}),$ (4)
where $\mathcal{F}(E_{m},E_{n})$ is the thermal occupation factor and $\mathcal{L}(E)=(\epsilon/\pi)/\left(E^{2}+\epsilon^{2}\right)$ the Lorentzian broadening, parameterized by $\epsilon$. Figures 3(a) and (b), with the transition lines normalized to their maximum value, $\widetilde{\mathcal{T}}=\mathcal{T}/\mathcal{T}_{\mathrm{max}}$, correspond to the system described by Figs. 2(a) and (c), respectively. While Fig. 3(a) which clearly shows the bulk bandgap, as we see a very limited transition probability around $E\approx 0$ and $\phi\approx\pi$, we can conclude that Fig. 3(b) displays the gap closing signature of topological superconductivity.
Unlike the majority of the microwave studies with a much smaller number of ABS Janvier et al. (2015); Wallraff et al. (2004); Tosi et al. (2019); van Woerkom et al. (2017); Hays et al. (2021) with theoretical results that we can recover by reducing the system size of our model JJs Bretheau et al. (2017); Dartiailh et al. (2021b); Elfeky et al. (2023b), the several hundred ABS considered in Fig. 3 complicate the identification of topological superconductivity. To address this, we resolve the calculated transition lines into pair and single-particle transitions, now replotted up to $f_{d}=40\;$GHz, compatible with relevant experimental setups for microwave detection, instead of the full ABS energy range of $2\Delta_{0}$ or $\sim 120\;$GHz.
By comparing these two types of transitions for the trivial [Figs. 3(e) and (f)] and topological phase [Figs. 3(g) and (h)], we see some striking differences. This is further illustrated in the inset of Fig. 3(g), which is the enlarged version of Fig. 2(c) focusing on low-energy ABS spectra near the topological gap closing and reopening, and the matching features in Fig. 3(g) and (h). The features marked by “A” and “B” in the inset denote the locations of topological gap closing and reopening and the MBS at $E\approx 0$ between these points. The matching V-shaped features seen in Fig. 3(g), also marked by “A” and “B,” resolve the pair transition lines around the gap closing and reopening, which therefore mark the microwave signatures of the topological superconductivity. On the other hand, the V-shaped single-particle transition lines seen in Fig. 3(h), marked by “C”, correspond to the transitions between the ABS levels in a region far from $E\approx 0$ state (marked by “C” in the inset) and therefore are not related to topological superconductivity. We also note that the features near zero frequency at $\phi\sim\pi/2,3\pi/2$ in Fig. 3(a), (b), (f) and (h) are related to single-particle transitions between ABS near $E\sim\Delta_{0}$, and are not topology related.
While, to the best of our knowledge, these here identified microwave signatures of topological superconductivity were not previously discussed, it is reassuring that for the trivial phase, we recover the behavior known from the previous JJ studies. For example, by reducing the number of ABS that we consider, we retain all the main features of the previous work on the weak-link JJs Tosi et al. (2019). This also includes the observed band gap around $\phi=\pi$ in Figs. 3(e) and (f), as expected for the trivial phase.
Figure 4: (a)-(c) The transition lines and (d)-(f) $\delta I(f,\phi)$ for the ABS from Fig. 2(c) with suppressed higher-energy single-particle transitions. The lines $\widetilde{\mathcal{T}}(f,\phi)$ in (a) are resolved into (b) pair and (c) single-particle transitions. The remaining transition lines in (c) are due to the energy-split overlapping MBS. (d) $\delta I(f,\phi)$ displays the gap closing and reopening. $\delta I(f,\phi)$ shown from (e) pair and (f) single-particle transitions, reveals the gap closing and opening as well as the transitions from $E\approx 0$ split MBS.
To provide further guidance to experimental work and previous $T=0$ results for a small number of Andreev bound states  Blais et al. (2021); Bretheau et al. (2014); Olivares et al. (2014), we consider the effects of the microwave drive on the JJ’s CPR, which affects the $f_{r}$ as discussed above. Without the microwave drive, the free energy of the JJ is $F=-k_{B}T\sum_{E_{n}}\mathrm{ln}\left[2\,\mathrm{cosh}\left(E_{n}/2k_{B}T\right)\right]$ Tinkham (1996), where $k_{B}$ is the Boltzmann constant and $\{E_{n}\}$ is the spectrum for the JJ. The CPR, $I(\phi)=(2e/\hbar)\,\partial F\,/\,\partial\phi$, is then
$I(\phi)$ $=-\frac{e}{\hbar}\,\sum_{E_{n}}\,\frac{\partial E_{n}}{\partial\phi}\,\mathrm{tanh}\left(\frac{E_{n}}{2k_{B}T}\right).$ (5)
Equation (5) shows the contribution of each energy level to the supercurrent, $(e/\hbar)\,\partial E_{n}/\partial\phi$, weighted by a thermal occupation $\mathrm{tanh}(E_{n}/2k_{B}T)$. We consider a change to $I(\phi)$ in Eq. (5) as a function of $f_{d}$, due to the microwave drive. At short timescales, we phenomenologically express the change in $I(\phi)$ as
$\delta I(f,\phi)$ $=\Gamma\,\sum_{m,n}\,\left(\frac{\partial E_{m}}{\partial\phi}-\frac{\partial E_{n}}{\partial\phi}\right)\,|\mathcal{M}_{mn}|^{2}$
×$\quad\times\,\mathcal{F}(E_{m},E_{n})\,\mathcal{L}(E_{mn}-hf_{d}),$ (6)
We choose a special form of the thermal occupation $\mathcal{F}(E_{m},E_{n})=f_{\mathrm{FD}}(E_{n})\,\left(1-f_{\mathrm{FD}}(E_{m})\right),$ where $f_{\mathrm{FD}}$ is the Fermi-Dirac function, to achieve the effect of suppressing single-particle transitions between the levels away from $E\approx 0$, due to the doubling of the fermionic degrees of freedom in the BdG Hamiltonian. This form of $\mathcal{F}(E_{m},E_{n})$ allows us to simulate the suppression of quasiparticle transitions between energy levels $E_{m,n}$ away from 0. The identification of topological signatures is simplified in the resulting transition lines $\widetilde{\mathcal{T}}$ and the shift in the current $\delta I(f,\phi)$. A counterpart plot with a conventional thermal occupation factor is discussed in SM SM_ . In the limit of $T=0\;$K and $\epsilon=0$, our Eq. (6) recovers the same form as in the published literature Blais et al. (2021); Bretheau et al. (2014); Olivares et al. (2014). While Eq. (6) does not include the $\phi$-dependent change of the slope and curvature of the ABS, our key results and trends are preserved by this simplification Park et al. (2020); SM_ .
For the JJ from Fig. 2(c), with single-particle transitions where $|E|\gtrsim k_{B}T$, we use this approach to plot the transition lines in Figs. 4(a)-(c) and $\delta I(f,\phi)$ of Eq. (6) in Figs. 4(d)-(f). Both the transition lines and $\delta I(f,\phi)$ reveal the signatures of the gap closing and opening, even for a system with hundreds of ABS. While these low-frequency transitions near the gap closing points appear in Figs. 4(b) and (e) for the pair transition, an even more striking feature becomes visible in Figs. 4(c) and (f) for the single-particle transitions. The suppression of the higher-energy single-particle transitions, including the feature “C” in Fig. 3(h), removes the background that competes with the signal from transitions between the energy-split MBS and higher energies, now observed prominently in Fig. 4(c) as a transition line between $\phi\sim 0.8\pi$ and $\phi\sim 1.2\pi$ with half of the slope of the V shapes marked by “A” and “C” in Fig. 3(g) Väyrynen et al. (2015); Murthy et al. (2020). This “half-slope” feature is also observed in Figs. 36 in the panels that depict topological transitions. The time-resolved capabilities of microwave spectroscopy could then also provide a detection of the non-Abelian statistics in planar JJs Zhou et al. (2020, 2022), which is supported by experimentally measured parameters, but is very challenging using conventional transport measurements.
Our overall features in Fig. 4 are consistent with other work on ABS in quantum dots Deacon et al. (2010); Zhang et al. (2022), constrictions Janvier et al. (2015); Wallraff et al. (2004) and nanowires Tosi et al. (2019); van Woerkom et al. (2017); Hays et al. (2021). This is important for our effort to expand the microwave spectroscopy as a sensitive probe for topological superconductivity in systems with a large number of ABS, where additional signatures become more pronounced by resolving pair and single-particle transitions. For example, one could apply our proposed detection to topological JJs based on 2D materials Xie et al. (2023), or systems with a dominant cubic SOC where the topological superconductivity is expected to be $f$-wave Alidoust et al. (2021); Luethi et al. (2023). It would be interesting to test if the enhanced signatures, obtained by removing the single-particle transitions (violating fermion parity conservation), could be implemented experimentally by identification and dynamical selection of parity in JJs, demonstrated in atomic contacts van Woerkom et al. (2017); Bretheau et al. (2013b) and weak links Ackermann et al. (2023); Wesdorp et al. (2024).
Figure 5: (a) ABS spectrum as a function of $B$ showing the lowest 240 levels with $\mu=1$ meV and $\phi=0.75\pi$. The gap closing and reopening, and the appearance of the zero-energy MBS at $B=0.24\;$T match the topological phase transition (see SM SM_ ). (b) Transition lines $\widetilde{\mathcal{T}}$ as a function of $B$. V feature indicating that the gap is closing and reopening is visible at $B=0.24\;$T near $f_{d}\sim 0$.
The V feature of topological gap closing and reopening is also seen in Fig. 5 from our numerical results for an experimental system where the JJ is phase biased and the Zeeman field is varied. We demonstrate the case where the phase bias is $\phi=0.75\pi$ (at $\phi=\pi$ the transition would be very close to $B=0\;$T). The ABS spectrum in Fig. 5(a) and the topological phase diagram in SM SM_ indicate a transition at $B=0.24\;$T. A corresponding V feature in the transition lines is observed in Fig. 5(b). The resolved pair transitions, and the $\delta\,I(f,B)$ plots in SM SM_ also demonstrate the V feature and the gap reopening, indicating that topological transitions may be observed in phase-biased microwave experiments.
Figure 6: (a) ABS spectrum as a function of $\phi$ showing the lowest 240 levels with $\mu=5$ meV and $B=0.5\;$T, for $L=4\;\mu$m. (b) Transition lines $\widetilde{\mathcal{T}}$ as a function of $B$. The double V feature, indicating the gap closing and reopening, is less sharp but visible around $\phi=0.75\pi$ and $\phi=1.25\pi$ near $f_{d}\sim 0$.
We next examine the influence of a higher chemical potential in a JJ with many ABS levels. The corresponding calculated results are shown in Fig. 6. By comparing the results for the spectra in Figs. 2(c) and 6(a), we see that at a higher $\mu$ will be less defined at a constant $B$ as a function of $\phi$, since a wider range of ABS levels will cross $E=0$. This will result in a reduced topological gap and less protected MBS, and will lead to a less sharp V-shaped signature in the transition lines [Fig. 6(b)]. A larger $B$ will shift the gap closing points further apart by shifting the ABS lines in the spectra, making the gap reopening signature more distinct. Similarly, a larger finite $L$ will open the topological gap further, having the same effect of making the signatures more distinct (see SM SM_ ). In our simulations with $\mu=5\;$meV and $L=4\;\mu$m, $\alpha=10\;$meV nm at $B=0.5\;$T, the gap reopening signature remains prominently visible but less sharp in the transition lines.
Our findings have various implications beyond topological superconductivity. Microwave spectroscopy of ABS could provide a sensitive probe of nonreciprocal phenomena in JJs Amundsen et al. (2024), while its temporal resolution could elucidate novel driving mechanism with time-dependent gate-control of SOC Monroe et al. (2022). As a result, even in the absence of any bias current in the JJ, it is possible to consider enhanced opportunities for superconducting electronics and neuromorphic computing using voltage pulses corresponding to the single flux quantum Tafuri (2019) (ed.); Holmes et al. (2013); Monroe et al. (2022); Likharev and Semenov (1991); Crotty et al. (2010). With decades of transport studies in junctions with only a single superconductor Kashiwaya and Tanaka (2000), microwave detection also offers a complementary approach to revisit various manifestations of ABS and unconventional superconductivity.
Note added–Recently, we were made aware of a publication Trif et al. (2018) titled “Current susceptibility as a probe of Majorana bound states in nanowire-based Josephson junctions,” which is relevant to our work.
Acknowledgements.
We thank Prof. Leonid Glazman for valuable discussions. This work is supported by the U.S. ONR MURI Award No. N000142212764 and U.S. ONR Award No. N000141712793. Computational resources were provided by the UB Center for Computational Research.